The following sample observations were randomly selected. 4 4 X. 6 3 15 13 11 95 13 15 7 Determine the correlation coefficient and interpret relationship between X and Y. the
Name Student ID Date Classwork of chapter21 1. A goatherd has the cost function cly)-Sy' where y is the number of tubs of goat cheese she makes per month. She faces a competitive market for goat cheese, with a price of $100 a tub. How many tubs should she produce per month? a) the square root of 100 (b) 25 (c) 10 (d) the square root of 20 (e) 5 2. A firm has a short run cost function c(y)-3y+ 11 for y> 0 and c(o)-7. The firm's quasi- fixed costs are: (a) 7 (c) 4 (d) 7.50. (e) impossible to determine from this information. 3. A firm has the short run total cost function c(y)-9y2 + 441 At what quantity of output is short run average cost minimized? (a) 7 (b) 3 (c) 49 (d) 0.43 (e) None of the above. 4. Mr. Dent Carr's total costs are 2s?+75s+ 100. If he repairs 25 cars, his average variable costs will be: () (a) 125 b) 129. (c) 175. (d) 250 (e) 87.50
Answer #1

5 875 3.5 IS 6as 61. 125 12 6562S 3875 3.S 156S 千 -0.8#5 0,チ65625 0-765625- 4- -0,875 6 I 105 1 265 6S 31S 9: 7 65625 2 9 = 3の2-16 = 13- 14 ว่า-1o.2s -チ5 4.s 0.93g .7S X and y asse 18.75 42 Juee, ot but git The .pa踧ano, deueu.如 (b) 오소 y 441 3a) (as) +15 50 +75= 135 Ave- =